# Combining inherent knowledge of vision-language models with unsupervised domain adaptation through self-knowledge distillation

### v2

@@ Abstract
Unsupervised domain adaptation (UDA) tries to overcome the tedious work of labeling data by leveraging a labeled source dataset and transferring its knowledge to a similar but different target dataset. On the other hand, current vision-language models exhibit astonishing zero-shot prediction capabilities. In this work, we combine knowledge gained through UDA with the inherent knowledge of vision-language models. In a first step, we generate the zero-shot predictions of the source and target dataset using the vision-language model. Since zero-shot predictions usually exhibit a large entropy, meaning that the class probabilities are rather evenly distributed, we first adjust the distribution to accentuate the winning probabilities. This is done using both source and target data to keep the relative confidence between source and target data. We then employ a conventional DA method, to gain the knowledge from the source dataset, in combination with self-knowledge distillation, to maintain the inherent knowledge of the vision-language model. We further combine our method with a gradual source domain expansion strategy (GSDE) and show that this strategy can also benefit by including zero-shot predictions. We conduct experiments and ablation studies on three benchmarks (OfficeHome, VisDA, and DomainNet) and outperform state-of-the-art methods. We further show in ablation studies the contributions of different parts of our algorithm.
@@ 1 Introduction
Deep neural networks have advanced the field of computer vision largely. However large amount of labeled data is required to train these networks. One method to overcome the need for tedious labeling is unsupervised domain adaptation, where knowledge from a similar but distinct labeled source dataset is transferred to an unlabeled target dataset.
On the other hand, vision-language models exhibit remarkable zero-shot prediction accuracy, even without any task-specific training data. It may even seem that through generating larger foundation models the need for domain adaptation fades. For example, on the DomainNet dataset, current vision-language models outperform state-of-the-art domain adaptation methods while only using zero-shot predictions.
@@
To maintain the inherent knowledge of the vision-language model we employ a knowledge distillation loss using the zero-shot predictions. However, since the zero-shot predictions usually exhibit a large entropy, meaning that the class probabilities are rather evenly distributed, we first adjust the distribution to accentuate the winning probability. To keep the relative confidence of the zero-shot prediction between source and target data we do this step for source and target data in a combined way. We use the knowledge distillation loss, to maintain the inherent knowledge, in combination with a conventional unsupervised domain adaptation method to benefit from the knowledge of the source dataset. As can be seen in Fig. 1 combining the two knowledge sources significantly increases the performance. We further combine our method with a gradual source domain expansion strategy (GSDE) [30] and show that this strategy can also benefit by including zero-shot predictions.
@@
  • We introduce self-knowledge distillation to maintain inherent knowledge of vision-language models using recalibrated zero-shot predictions and combine them with a conventional UDA method.
  • We further employ a gradual source domain expansion GSDE) and enhance the strategy by introducing zero-shot predictions into it.
@@
  • We show that the method is applicable for both CNN-based and transformer-based backbones.
Figure 2: Process flow of our algorithm. In a first step, the zero-shot predictions of source $\mathcal{D}_{s}$ and target $\mathcal{D}_{t}$ dataset are estimated. The zero-shot distributions are calibrated jointly to accentuate the winning probability. The network is trained using a classification loss $L_{CE}$ for the source data, a knowledge distillation loss employing the shifted zero-shot predictions $\tilde{y}_{o}$, and an adversarial adaptation loss $L_{DA}$. We further employ strongly augmented versions of source and target data as input. Since we do not differentiate between weakly and strongly augmented versions for the losses, their paths are combined after the visual encoder.
@@ 2.1 Unsupervised domain adaptation
The goal of unsupervised domain adaptation is to overcome the domain shift between the source and the target dataset. Surveys for this task can be found in [41], [38]. One approach is to align the feature space distributions between the two datasets. Domain-adversarial neural network (DANN) [5] introduces an adversarial approach to achieve this. A domain classifier is trained to distinguish the feature space of source and target samples. The work introduced a gradient reversal layer. This layer inverts the gradients and thereby inverts the training objective, meaning that the backbone is trained to generate features that are indistinguishable for the domain classifier, generating so-called domain invariant features. BIWAA introduces a feature re-weighting approach based on their contribution to the classifier. CDAN [15] extended DANN by multilinear conditioning the domain classifier with the classifier predictions. Moving semantic transfer network [32] introduced a moving semantic loss.
Another approach employs information maximization or entropy minimization to generate more accurate target predictions. SHOT [14] exploits both information maximization and self-supervised pseudo-labeling to increase the performance of the target domain. SENTRY [21] uses prediction consistency among different augmentations of an image to either minimize its entropy or maximize it in case of inconsistent predictions.
Following the rise of vision transformers, approaches specially tailored for these backbones have gained popularity. TVT [34] introduced a Transferability Adaption Module that employs a patch-level domain discriminator to measure the transferability of patch tokens, and injects learned transferability into the multi-head self- attention block of a transformer. CDTrans [33] introduced a triple-branch transformer framework with cross-attention between source and target domains. PMTrans [40] introduced PatchMix which builds an intermediate domain by mixing patches of source and target images. The mixing process uses a learnable Beta distribution and attention-based scoring function to assign label weights to each patch.
@@ 2.3 Domain adaptation for vision-language models
Few approaches make use of prompt learning [39]. This approach trains the prompt’s context words for the text encoder with learnable vectors while the encoder weights are frozen. DAPL [7] employs a prompt learning approach to learn domain-agnostic context variables and domain-specific context variables for unsupervised domain adaptation. This approach freezes the text and vision encoder during training. AD-CLIP [25] introduces a prompt learning scheme that entirely leverages the visual encoder of CLIP and introduces a small set of learnable projector networks. MPA [2] takes a similar approach of prompt learning for multi-source unsupervised domain adaptation. In a first step, individual prompts are learned to minimize the domain gap through a contrastive loss. Then, MPA employs an auto-encoding process and aligns the learned prompts by maximizing the agreement of all the reconstructed prompts. DALL-V [35] employs large language-vision models for the task of source-free video domain adaptation. They distill the world prior and complementary source model information into a student network tailored for the target. This approach also freezes the vision encoder and only learns an adapter on top of the vision encoder.
Unlike all of the above methods, we do not freeze the vision encoder during the training. Particularly, we employ the zero-shot predictions and use them in a knowledge distillation step to retain the knowledge of the language-vision model. Furthermore, we combine our method with a conventional UDA approach.
Table 1: Accuracy results on Office-Home dataset. The best results are displayed in bold and the runner-up results are underlined. Methods using a ResNet-50 backbone are on top, and methods using a transformer-based backbone are on the bottom. ZS indicates zero-shot results of CLIP.
@@
ZS [22] 48.5 81.2 83.5 72.6 81.2 83.5 72.6 48.5 83.5 72.6 48.5 81.2 71.5
SKD (Ours) 61.6 86.8 86.7 78.0 87.4 86.8 77.3 61.0 87.1 79.6 64.1 88.9 78.8
@@
ZS [22] 66.8 89.1 89.7 83.3 89.1 89.7 83.3 66.8 89.7 83.3 66.8 89.1 82.2
SKD (Ours) 79.6 93.7 92.7 89.5 93.7 92.9 89.1 81.1 92.6 90.2 81.6 94.2 89.2
@@ 3 Methodology
In unsupervised domain adaptation, the goal is to transfer knowledge from a source dataset $\mathcal{D}_{s}$ consisting of $n_{s}$ labeled images $\mathcal{D}_{s}={(x_{i,s},y_{i,s})}_{i=1}^{n_{s}}$ to an unlabeled target dataset $\mathcal{D}_{t}$ containing $n_{t}$ unlabeled data $\mathcal{D}_{t}={(x_{i,t})}_{i=1}^{n_{t}}$. In this work we focus on combining the inherent knowledge of vision-language models with the knowledge gained from the knowledge transfer of the source domain, meaning we further have access to the zero-shot predictions $y_{o}$ for each sample.
For the knowledge transfer from the source to the target domain, we employ the widely established adversarial domain adaptation method CDAN as it was shown to be effective for both convolutional- and transformer-based network architectures [11].
To preserve the inherent knowledge of the vision-language model we employ a knowledge distillation loss [8] using the zero-shot predictions. However, since the zero-shot predictions exhibit a relatively wide spread of probabilities among the classes (see Fig. 3), we first adjust the distribution to accentuate the winning probabilities, so that they are more informative for the knowledge distillation loss. The process flow of our algorithm can be seen in Fig. 2
We further employ data augmentation and a gradual source domain expansion strategy to further boost the performance.
For a single run, we use the three losses:
@@
3.1 Knowledge distillation loss
Vision-language models already achieve an incredible zero-shot accuracy for unseen data. To preserve this knowledge we employ a knowledge distillation loss. However, since the predictions are rather evenly spread among all classes, we adjust the distribution (from $y_{o}$ to $\tilde{y}_{o}=\alpha\cdot y_{o}$) before employing them in the KD-loss (see Fig. 3). In a normal KD setting, the teacher’s winning prediction score usually tends to approach a one-hot vector. The goal is to mimic this distribution while maintaining the relative prediction probabilities. To accomplish that we estimate a factor $\alpha$ so that the mean winning probability equals to $\tau$. We employ a value of $\tau=0.9$ in our experiments. In the estimation, we equally factor the source and target domain to further maintain the relative prediction confidence between the two domains.
$\frac{1}{2n_{s}}\sum_{n_{s}}\max(\sigma(y_{o,i,s}\cdot\alpha))+\frac{1}{2n_{t}}\sum_{n_{t}}\max(\sigma(y_{o,i,t}\cdot\alpha))=\tau$ (2)
We then employ the shifted zero-shot predictions for the knowledge distillation loss:
@@ 3.1 Knowledge distillation loss
Figure 3: Distribution of winning probabilities of zero-shot predictions $y_{o}$ and their adjusted distribution $\tilde{y}_{o}$ for the clipart and real domain of OfficeHome. While the zero-shot predictions have low winning probabilities, the adjusted distribution resembles much more one-hot vectors, while maintaining their relative probabilities, as can be seen from the long tail of low winning probabilities.
Table 2: Accuracy results on VisDA dataset. The best results are displayed in bold and the runner-up results are underlined. Methods using a ResNet-101 backbone are on top, and methods using a transformer-based backbone are on the bottom. ZS indicates zero-shot results of CLIP.
Method plane bcycl bus car horse knife mcycl person plant sktbrd train truck Avg
@@
ZS [22] 98.0 83.3 90.8 65.9 97.8 83.0 93.4 67.7 85.2 90.0 94.9 66.4 84.7
SKD (Ours) 98.8 87.8 92.0 72.0 98.7 93.4 94.8 75.1 92.5 96.0 95.9 72.0 89.1
TVT [34] 82.9 85.6 77.5 60.5 93.6 98.2 89.4 76.4 93.6 92.0 91.7 55.7 83.1
SHOT [14] 99.5 91.8 88.7 65.1 98.6 98.0 96.0 66.1 95.1 98.9 96.8 52.4 87.3
@@
ZS [22] 99.3 91.5 92.9 71.0 99.4 92.7 94.4 75.6 84.7 97.4 95.5 69.0 88.6
SKD (Ours) 99.6 95.2 94.3 74.5 99.6 98.0 95.9 79.8 89.8 99.2 96.6 71.8 91.2
3.2 Adversarial loss
We employ conditional adversarial domain adaptation loss CDAN:
$L_{AD}=L_{BCE}(G_{d}((f_{i}\otimes p_{i}),d_{i}))$ (4)
where $G_{d}$ is the domain classification network. $f_{i}$ are the features of sample $i$, $p_{i}$ the class probabilities and $d_{i}$ the domain label. CDAN employs a gradient reversal layer to inverse the training objective. This means that while the domain classifier is trained to distinguish the domain (source of target) of the respective sample, the objective of the feature extractor is to extract features that are indistinguishable for the domain classifier, thereby generating domain invariant features.
3.3 Gradual source domain expansion
We further employ the gradual source domain expansion strategy [30]. This strategy trains the network iteratively $N$ times, each time resetting the network weights. However, each new run $n$ employs the $\frac{n-1}{N}$ highest scoring target samples with their respective pseudo-labels estimated from the previous run as additional source data. To account for the additional knowledge available from the zero-shot predictions, we slightly modify the algorithm. While the original algorithm only considers the final predictions of the previous run to choose the pseudo-source data, we combine the final predictions with the zero-shot predictions:
$s(\hat{y})=p(y_{n-1})+\frac{1}{2^{n}}p(\tilde{y}_{o})$ (5)
where $s(\hat{y})$ is the scores of the samples for being chosen as pseudo-source samples, and $p(y_{n-1})$ is the final predictions of the previous run. As the performance increases over the runs, we decrease the influence of the zero-shot predictions with every run. We use a value of $N=5$ in our experiments.
3.4 Batch norm layer adjustment

### v4

# Combining inherent knowledge of vision-language models with unsupervised domain adaptation through strong-weak guidance

@@ Abstract
Unsupervised domain adaptation (UDA) tries to overcome the tedious work of labeling data by leveraging a labeled source dataset and transferring its knowledge to a similar but different target dataset. Meanwhile, current vision-language models exhibit remarkable zero-shot prediction capabilities. In this work, we combine knowledge gained through UDA with the inherent knowledge of vision-language models. We introduce a strong-weak guidance learning scheme that employs zero-shot predictions to help align the source and target dataset. For the strong guidance, we expand the source dataset with the most confident samples of the target dataset. Additionally, we employ a knowledge distillation loss as weak guidance. The strong guidance uses hard labels but is only applied to the most confident predictions from the target dataset. Conversely, the weak guidance is employed to the whole dataset but uses soft labels. The weak guidance is implemented as a knowledge distillation loss with (adjusted) zero-shot predictions. We show that our method complements and benefits from prompt adaptation techniques for vision-language models. We conduct experiments and ablation studies on three benchmarks (OfficeHome, VisDA, and DomainNet), outperforming state-of-the-art methods. Our ablation studies further demonstrate the contributions of different components of our algorithm.
@@ 1 Introduction
Deep neural networks have significantly advanced the field of computer vision. However, training these networks requires a large amount of labeled data. Unsupervised domain adaptation (UDA) offers a solution by transferring knowledge from a labeled similar but distinct source dataset to an unlabeled target dataset, reducing the need for extensive labeling.
On the other hand, vision-language models exhibit remarkable zero-shot prediction accuracy, even without any task-specific training data. For instance, on the DomainNet dataset, current vision-language models’ zero-shot predictions outperform state-of-the-art domain adaptation without the need for a source dataset. This raises the question of the importance of domain adaption method give the rapid emergence of larger foundation models.
@@
To preserve the inherent knowledge of the vision-language model we employ a strong-weak guidance. The strong guidance is applied only for the most confident samples using the hard predictions (i.e. pseudo-label) in form of a source dataset expansion. Conversely, the weak guidance employs soft predictions and is applied to all samples in form of a knowledge distillation loss. We show that adjusting the output probabilities by accentuating the winning probabilities enhances the effectiveness of knowledge distillation.
By combining our strong-weak guidance with a conventional unsupervised domain adaptation method we can significantly increase the performance as can be seen in Fig. 1. This highlights the potential of combining inherent knowledge of vision-language models with knowledge transferred from a source domain.
While much research on domain adaptation with CLIP models has focused on adapting the text prompt, our method focuses on adapting the visual encoder, while maintaining its inherent knowledge. We show that these two approaches are complementary, and show that our method benefits from combining it with the prompt adaptation method DAPL [7].
@@
  • We introduce a strong-weak guidance scheme to exploit inherent knowledge of vision-language models and combine it with a conventional UDA method.
  • We introduce a combination of weak guidance based on knowledge distillation of adjusted zero-shot predictions and strong guidance based on source dataset expansion.
@@
  • We show that the method is applicable for both CNN-based and ViT-based backbones and can be combined with other UDA methods.
@@ 2.1 Unsupervised domain adaptation
The goal of unsupervised domain adaptation is to overcome the domain shift between the source and the target dataset. Surveys for this task can be found in [40], [37]. One approach is to align the feature space distributions between the two datasets. Domain-adversarial neural network (DANN) [5] introduces an adversarial approach to achieve this. A domain classifier is trained to distinguish the feature space of source and target samples. The work introduced a gradient reversal layer. This layer inverts the gradients and thereby inverts the training objective, meaning that the backbone is trained to generate features that are indistinguishable for the domain classifier, generating so-called domain invariant features. BIWAA [29] introduces a feature re-weighting approach based on their contribution to the classifier. CDAN [17] extended DANN by multilinear conditioning the domain classifier with the classifier predictions. Moving semantic transfer network [31] introduced a moving semantic loss.
Another approach aims for the minimization of joint-error [36]. Others approaches employ information maximization or entropy minimization to generate more accurate target predictions. SHOT [16] exploits both information maximization and self-supervised pseudo-labeling to increase the performance of the target domain. SENTRY [23] uses prediction consistency among different augmentations of an image to either minimize its entropy or maximize it in case of inconsistent predictions.
Following the rise of vision transformers approaches specially tailored for these backbones have gained popularity. TVT [33] introduced a Transferability Adaption Module that employs a patch-level domain discriminator to measure the transferability of patch tokens, and injects learned transferability into the multi-head self- attention block of a transformer. CDTrans [32] introduced a triple-branch transformer framework with cross-attention between source and target domains. PMTrans [39] introduced PatchMix which builds an intermediate domain by mixing patches of source and target images. The mixing process uses a learnable Beta distribution and attention-based scoring function to assign label weights to each patch.
@@ 2.2 Vision-language models
Figure 2: Process flow of our algorithm. In a first step, the zero-shot predictions of source $\mathcal{D}_{s}$ and target $\mathcal{D}_{t}$ dataset are estimated. We shift the zero-shot predictions distribution through a temperature parameter in the softmax to accentuate the winning probability. Based on the zero-shot predictions, we extend the source dataset with high confident target samples. These samples are treated as source data, with their respective pseudo-labels and represent the strong guidance of our method. The network is then trained using a classification loss $L_{CE}$ for the (expanded) source data, a knowledge distillation loss $L_{KD}$ employing the shifted zero-shot predictions $\tilde{y}_{o}$, and an adversarial adaptation loss $L_{DA}$. The knowledge distillation loss represents the weak guidance of the method, as it is employed for all samples and uses the soft zero-shot predictions.
@@ 2.3 Domain adaptation for vision-language models
Few approaches make use of prompt learning [38]. This approach trains the prompt’s context words for the text encoder with learnable vectors while the encoder weights are frozen. DAPL [7] employs a prompt learning approach to learn domain-agnostic context variables and domain-specific context variables for unsupervised domain adaptation. This approach freezes the text and vision encoder during training. AD-CLIP [26] introduces a prompt learning scheme that entirely leverages the visual encoder of CLIP and introduces a small set of learnable projector networks. PADCLIP [12] aims for a debiasing zero-shot predictions and employs a self-training paradigm, we employ adjusted prediction scores instead of hard pseudo-labels and use an adversarial approach to overcome the domain gap. EUDA [13] employs prompt task-dependant tuning and a visual feature refinement module in combination with pseudo-labeling from semi-supervised classifiers. MPA [3] takes a similar approach of prompt learning for multi-source unsupervised domain adaptation. In a first step, individual prompts are learned to minimize the domain gap through a contrastive loss. Then, MPA employs an auto-encoding process and aligns the learned prompts by maximizing the agreement of all the reconstructed prompts. DALL-V [34] employs large language-vision models for the task of source-free video domain adaptation. They distill the world prior and complementary source model information into a student network tailored for the target. This approach also freezes the vision encoder and only learns an adapter on top of the vision encoder.
Unlike most of the above methods, we do not freeze the vision encoder during the training. While most of the mentioned methods focus on adapting the text prompts, we focus on adapting the visual extractor while retaining the inherent knowledge through our strong-weak guidance. We also show our method is complementary to and benefits from prompt learning methods.
Table 1: Accuracy results on Office-Home dataset. The best results are displayed in bold and the runner-up results are underlined. Methods using a ResNet-50 backbone are on top, and methods using a transformer-based backbone are on the bottom. CLIP indicates zero-shot results of CLIP, and methods based on it are listed below it. Our methods are on the bottom, DAPL marks the combination with the prompt learning method. V1 stands for version 1 and V2 for version 2 respectively.
@@
DAPL [7] 54.1 84.3 84.8 74.4 83.7 85.0 74.5 54.6 84.8 75.2 54.7 83.8 74.5
AD-CLIP [26] 55.4 85.2 85.6 76.1 85.8 86.2 76.7 56.1 85.4 76.8 56.1 85.5 75.9
PADCLIP [12] 57.5 84.0 83.8 77.8 85.5 84.7 76.3 59.2 85.4 78.1 60.2 86.7 76.6
EUDA [13] 58.1 85.0 84.5 77.4 85.0 84.7 76.5 58.8 85.7 75.9 60.4 86.4 76.5
SWG V1 64.3 86.2 86.8 78.8 87.6 86.6 79.4 64.5 86.9 80.9 66.1 88.9 79.7
SWG V2 66.6 87.3 87.0 79.6 87.4 86.7 78.8 66.5 87.2 80.4 68.0 88.8 80.3
SWG DAPL V1 70.8 89.8 88.9 81.0 89.9 89.0 81.2 70.8 89.3 84.0 74.0 92.2 83.4
SWG DAPL V2 73.8 89.8 89.3 81.7 90.0 89.0 81.6 74.4 89.3 84.7 76.4 91.9 84.3
[…]
@@
DAPL [7] 70.6 90.2 91.0 84.9 89.2 90.9 84.8 70.5 90.6 84.8 70.1 90.8 84.0
AD-CLIP [26] 70.9 92.5 92.1 85.4 92.4 92.5 86.7 74.3 93.0 86.9 72.6 93.8 86.1
PADCLIP [12] 76.4 90.6 90.8 86.7 92.3 92.0 86.0 74.5 91.5 86.9 79.1 93.1 86.7
EUDA [13] 78.2 90.4 91.0 87.5 91.9 92.3 86.7 79.7 90.9 86.4 79.4 93.5 87.3
SWG V1 81.5 93.5 92.7 89.5 93.3 92.8 89.2 82.4 92.6 90.2 82.7 94.3 89.6
SWG V2 81.6 93.6 92.7 89.6 93.3 93.1 89.1 83.1 92.9 90.4 83.1 94.1 89.7
SWG DAPL V1 85.8 93.8 93.5 89.4 94.0 93.9 90.0 86.8 94.2 91.6 87.4 95.3 91.3
SWG DAPL V2 87.1 93.9 94.1 91.1 94.3 93.8 91.1 89.0 94.2 92.3 89.4 95.8 92.2
@@ 3 Methodology
In unsupervised domain adaptation, the goal is to transfer knowledge from a source dataset $\mathcal{D}_{s}$ consisting of $n_{s}$ labeled images $\mathcal{D}_{s}={(x_{i,s},y_{i,s})}_{i=1}^{n_{s}}$ to an unlabeled target dataset $D_{t}$ containing $n_{t}$ unlabeled data $\mathcal{D}_{t}={(x_{i,t})}_{i=1}^{n_{t}}$. In this work we focus on combining the inherent knowledge of vision-language models with the knowledge gained from the knowledge transfer of the source domain, meaning we further have access to the zero-shot predictions $y_{o}$ for each sample.
For the knowledge transfer from the source to the target domain, we employ the widely established adversarial domain adaptation method CDAN [17] as it was shown to be effective for both convolutional- and transformer-based network architectures [11].
To preserve the inherent knowledge of the vision-language model we employ a strong-weak guidance scheme. For the strong guidance, a hard pseudo-label is employed, but only for the most confident samples. We employ a source domain expansion inspired by GSDE [30] that copies the most confident target samples into the source domain with their predicted pseudo-label.
For the weak guidance, we employ a knowledge distillation loss [8] using the zero-shot predictions. However, since the zero-shot predictions exhibit a relatively wide spread of probabilities among the classes, we first accentuate the winning probabilities. The process flow of our algorithm can be seen in Fig. 2.
We use the three losses:
@@
We also show that our method is complementary to prompt learning methods by combining it with Domain Adaptation via Prompt Learning (DAPL) [7].
3.1 Strong guidance - Source Domain Expansion:
For the strong guidance, we employ a source domain expansion strategy. The highest scoring zero-shot predictions are copied and appended to the source dataset with their respective pseudo-labels as pseudo-source data. We employ this source domain expansion strategy since it was shown to be more effective than simply adding a CE-loss for the respective target samples (as was shown in [30]). During training no distinction between source data and pseudo-source data is made.
The paper that inspired the strategy for our strong guidance employed the source domain expansion in an iterative procedure. The network was trained for several runs, each time with an increasing amount of pseudo-source data.
In this work, we experiment with 2 different versions. The first version trains the network a single time. The second version employs an iterative training scheme training for 2 runs. In the first run, the source domain expansion is done regarding the zero-shot predictions, and in the second run it is done regarding a mixture of zero-shot predictions and predictions generated from the previous run:
$s(\hat{y})=p(y_{n-1})+\frac{1}{2}p(\tilde{y}_{o})$ (2)
where $s(\hat{y})$ depicts the scores of the samples for being chosen as pseudo-source samples, $p(y_{n-1})$ depicts the predictions of the previous run, and $p(\tilde{y}_{o})$ depicts the zero-shot predictions.
While the first version trains for a single run, and therefore no additional computational time is required, the second version trains for two runs, which require double the training time.
3.2 Weak guidance - Knowledge distillation loss
For the weak guidance, we employ a knowledge distillation loss [8] using the zero-shot predictions. However, since the predictions are rather evenly spread among all classes, we accentuate the winning probabilities before employing them in the KD-loss (see Fig. 2). This is done through a temperature parameter $T$ in the softmax function. In a normal KD setting, the teacher’s winning prediction score usually tends to approach a one-hot vector. The goal is to mimic this setting while maintaining the relative prediction probabilities. To accomplish that we estimate the temperature parameter $T$ so that the mean winning probability over all source and target data equals to $\tau$. We employ a value of $\tau=0.9$ in our experiments. In the estimation, we equally factor the source and target domain to further maintain the relative prediction confidence between the two domains.
$\frac{1}{2n_{s}}\sum_{n_{s}}\max(\sigma(y_{o,i,s}/T))+\frac{1}{2n_{t}}\sum_{n_{t}}\max(\sigma(y_{o,i,t}/T))=\tau$ (3)
$y_{o,i,s}$ depicts logits of the $i$-th sample of the $s$ource data the $\sigma$ depicts the softmax-function. We then employ the adjusted zero-shot predictions for the knowledge distillation loss:
@@ 3.2 Weak guidance - Knowledge distillation loss
3.3 Adversarial loss
We employ conditional adversarial domain adaptation loss CDAN:
$L_{AD}=L_{BCE}(G_{d}((f_{i}\otimes p_{i}),d_{i}))$ (5)
where $G_{d}$ is the domain classification network. $f_{i}$ are the features of sample $i$, $p_{i}$ the class probabilities and $d_{i}$ the domain label. CDAN employs a gradient reversal layer to inverse the training objective. This means that while the domain classifier is trained to distinguish the domain (source or target) of the respective sample, the objective of the feature extractor is to extract features that are indistinguishable for the domain classifier, thereby generating domain invariant features.
Table 2: Accuracy results on VisDA dataset. The best results are displayed in bold and the runner-up results are underlined. Methods using a ResNet-101 backbone are on top, and methods using a transformer-based backbone are on the bottom. CLIP indicates zero-shot results of CLIP, and methods based on it are listed below it. Our methods are on the bottom, DAPL marks the combination with the prompt learning method. V1 stands for version 1 and V2 for version 2 respectively.
Method - RN101 plane bcycl bus car horse knife mcycl person plant sktbrd train truck Avg
@@ 3.3 Adversarial loss
DAPL [7] 97.8 83.1 88.8 77.9 97.4 91.5 94.2 79.7 88.6 89.3 92.5 62.0 86.9
AD-CLIP [26] 98.1 83.6 91.2 76.6 98.1 93.4 96.0 81.4 86.4 91.5 92.1 64.2 87.7
PADCLIP [12] 96.7 88.8 87.0 82.8 97.1 93.0 91.3 83.0 95.5 91.8 91.5 63.0 88.5
EUDA [13] 97.2 89.3 87.6 83.1 98.4 95.4 92.2 82.5 94.9 93.2 91.3 64.7 89.2
SWG V1 99.0 85.2 91.3 72.9 98.8 96.4 96.2 83.0 91.0 96.2 94.0 75.0 89.9
SWG V2 98.9 87.2 91.6 73.8 98.8 96.1 96.3 83.2 91.8 95.7 94.2 74.5 90.2
SWG DAPL V1 98.9 87.3 92.8 81.1 98.0 95.6 96.0 82.8 92.4 96.3 93.1 66.4 90.1
SWG DAPL V2 98.9 87.5 92.3 82.5 98.3 94.2 95.7 80.7 91.8 96.8 94.5 66.0 89.9
Method - ViT plane bcycl bus car horse knife mcycl person plant sktbrd train truck Avg
@@
DAPL [7] 99.2 92.5 93.3 75.4 98.6 92.8 95.2 82.5 89.3 96.5 95.1 63.5 89.5
AD-CLIP [26] 99.6 92.8 94.0 78.6 98.8 95.4 96.8 83.9 91.5 95.8 95.5 65.7 90.7
PADCLIP [12] 98.1 93.8 87.1 85.5 98.0 96.0 94.4 86.0 94.9 93.3 93.5 70.2 90.9
EUDA [13] 98.4 94.3 89.0 85.4 98.5 98.3 96.1 86.3 95.1 95.2 92.5 70.9 91.7
SWG V1 99.8 93.8 93.7 73.7 99.6 98.3 97.3 82.1 91.0 99.1 96.8 70.0 91.3
SWG V2 99.7 94.0 93.6 73.7 99.6 98.4 97.5 81.4 92.2 99.3 96.7 71.4 91.5
SWG DAPL V1 99.6 94.9 94.6 78.3 99.1 98.4 96.7 85.9 92.7 99.1 95.8 69.4 92.0
SWG DAPL V2 99.5 95.5 94.6 79.1 99.1 97.8 96.9 86.5 92.8 99.1 95.9 70.7 92.3
Table 3: Accuracy results on DomainNet dataset.
MCD clp inf pnt qdr rel skt Avg SCDA clp inf pnt qdr rel skt Avg CDTrans clp inf pnt qdr rel skt Avg
clp - 15.4 25.5 3.3 44.6 31.2 24.0 clp - 18.6 39.3 5.1 55.0 44.1 32.4 clp - 29.4 57.2 26.0 72.6 58.1 48.7
inf 24.1 - 24.0 1.6 35.2 19.7 20.9 inf 29.6 - 34.0 1.4 46.3 25.4 27.3 inf 57.0 - 54.4 12.8 69.5 48.4 48.4
pnt 31.1 14.8 - 1.7 48.1 22.8 23.7 pnt 44.1 19.0 - 2.6 56.2 42.0 32.8 pnt 62.9 27.4 - 15.8 72.1 53.9 46.4
qdr 8.5 2.1 4.6 - 7.9 7.1 6.0 qdr 30.0 4.9 15.0 - 25.4 19.8 19.0 qdr 44.6 8.9 29.0 - 42.6 28.5 30.7
rel 39.4 17.8 41.2 1.5 - 25.2 25.0 rel 54.0 22.5 51.9 2.3 - 42.5 34.6 rel 66.2 31.0 61.5 16.2 - 52.9 45.6
skt 37.3 12.6 27.2 4.1 34.5 - 23.1 skt 55.6 18.5 44.7 6.4 53.2 - 35.7 skt 69.0 29.6 59.0 27.2 72.5 - 51.5
Avg 28.1 12.5 24.5 2.4 34.1 21.2 20.5 Avg 42.6 16.7 37.0 3.6 47.2 34.8 30.3 Avg 59.9 25.3 52.2 19.6 65.9 48.4 45.2
SSRT clp inf pnt qdr rel skt Avg PMTrans clp inf pnt qdr rel skt Avg EUDA clp inf pnt qdr rel skt Avg
clp - 33.8 60.2 19.4 75.8 59.8 49.8 clp - 34.2 62.7 32.5 79.3 63.7 54.5 clp - 54.8 69.9 35.3 85.1 67.4 62.5
inf 55.5 - 54.0 9.0 68.2 44.7 46.3 inf 67.4 - 61.1 22.2 78.0 57.6 57.3 inf 70.2 - 68.5 16.6 82.2 65.9 60.7
pnt 61.7 28.5 - 8.4 71.4 55.2 45.0 pnt 69.7 33.5 - 23.9 79.8 61.2 53.6 pnt 72.4 54.6 - 29.5 83.0 66.4 61.2
qdr 42.5 8.8 24.2 - 37.6 33.6 29.3 qdr 54.6 17.4 38.9 - 49.5 41.0 40.3 qdr 73.1 50.8 64.3 - 81.2 62.3 66.3
rel 69.9 37.1 66.0 10.1 - 58.9 48.4 rel 74.1 35.3 70.0 25.4 - 61.1 53.2 rel 75.5 56.1 74.6 30.2 - 65.6 60.4
skt 70.6 32.8 62.2 21.7 73.2 - 52.1 skt 73.8 33.0 62.6 30.9 77.5 - 55.6 skt 74.9 56.2 70.2 32.3 80.3 - 62.8
Avg 60.0 28.2 53.3 13.7 65.3 50.4 45.2 Avg 67.9 30.7 59.1 27.0 72.8 56.9 52.4 Avg 73.2 54.5 69.5 28.8 82.4 65.5 62.3
PADCLIP clp inf pnt qdr rel skt Avg SWG V1 clp inf pnt qdr rel skt Avg SWG V2 clp inf pnt qdr rel skt Avg
clp - 55.1 71.1 36.8 84.2 68.1 63.1 clp - 58.4 73.1 25.7 85.0 71.9 62.8 clp - 58.2 74.0 30.4 85.4 72.7 64.2
inf 73.6 - 70.6 18.0 83.5 66.6 62.5 inf 78.7 - 73.1 24.7 84.8 70.8 66.4 inf 79.7 - 73.7 28.5 85.3 71.2 67.7
pnt 75.4 54.3 - 32.0 83.5 67.2 62.5 pnt 79.6 58.3 - 26.0 84.6 71.3 63.9 pnt 80.3 58.4 - 27.8 84.8 71.9 64.6
qdr 74.6 53.6 70.0 - 83.1 66.1 69.5 qdr 76.5 56.2 70.6 - 84.0 69.2 71.3 qdr 76.3 53.5 71.4 - 83.8 69.6 70.9
rel 76.4 54.9 72.7 31.7 - 67.5 60.6 rel 80.4 58.5 73.9 24.4 - 71.4 61.7 rel 81.2 58.8 74.4 29.6 - 72.0 63.2
skt 76.3 54.9 71.7 34.9 83.6 - 64.3 skt 80.6 58.1 73.6 25.1 84.8 - 64.5 skt 81.6 58.7 74.8 30.5 85.3 - 66.2
Avg 75.3 54.6 71.2 30.7 83.6 67.1 63.7 Avg 79.2 57.9 72.9 25.2 84.6 70.9 65.1 Avg 79.8 57.5 73.7 29.4 84.9 71.5 66.1
3.4 Further improvements:
3.4.1 Batch norm layer adjustment